Calculate the pH of a solution prepared by dissolving 0.30 g of Na2O in enough water to make 0.500 L of solution.
Answer #1

molar mass of Na2O= 23*2+16 =62 g/mole

moles of Na2O= mass/molar mass = 0.3/62=0.0048

volume of solution =0.5L

the dissolution of Na2O is Na2O+H2O ------>2NaOH

1 mole of Na2O gives rise to 2 moles of NaOH. hence moles of NaOH= 2*0.0048=0.0096

concentration of NaOH= 0.0096/0.5= 0.0192M

NaOH ionizes as NaOH-------->Na++OH-

[OH-]= 0.0192, pOH= -log(OH-)= 1.72

pH= 14-pOH= 14-1.72=12.28